i^2+i^4+i^6+i8+i10+i12+i14+i16+i18+i20+i22+i24+i26+i28.....+2n terms =

Grade:9

2 Answers

Arun
4 years ago
Dear student
 
it can be written as ( – 1 + 1) +( – 1 + 1) + ( – 1 + 1) + ( – 1 + 1) + …....n times
 
 = 0 + 0 + 0 +.... = 0
Vikas TU
4 years ago
Dear student 
i^2 = -1 
i^4 = 1 
So, Here you can put the values , 
i^6 = i^2 .
and i^8 = i^4 
Expand  similarly . 
The final ans would be 0. 
Hope this explnation helps .
Good Luck